OpenStudy (gabylovesyou):

Two more questions!!Consider the line that passes through the point (3, -2) and has a slope of 2. Part 1: Write the equation of this line using point-slope form. (2 points) Part 2: Using your equation from part 1, rewrite this equation in slope-intercept form. Make sure to show all of your work. (2 points) Part 3: Using your equation from part 2, rewrite this equation in standard form. Make sure to show all of your work. (2 points)

OpenStudy (anonymous):

http://lmgtfy.com/?q=point+slope+form

OpenStudy (ny,ny):

i dont know how to put it in point slope form, but i do know how to write the equation in slope intercept form. y=mx+b -2=2(3)+b -2=6+b subtract 6 from both sides... -8=b so the equation is y=2x-8 see what i did?

OpenStudy (anonymous):

point slope form is y - y1 = m ( x - x1) so (x1, y1) is the point (3,-2) and slope m=2

OpenStudy (anonymous):

point slope form y-(-2) = 2(x-3) slope intercept form y=2x-8 standard form 2x-y=8

OpenStudy (ny,ny):

@chenna you are right.

OpenStudy (ny,ny):

you can simplify the point slope form equation further.

OpenStudy (anonymous):

she will never learn if you keep giving her charity answers

OpenStudy (ny,ny):

awe, whatev.